OpenStudy (anonymous):

can someone help me please!!!!???? how would you write ' 12 less than a number times 2' into an algebraic translation?

OpenStudy (anonymous):

do you know how to designate an unknown number in algebraic writing?

OpenStudy (anonymous):

that's the first step to figuring it out.

OpenStudy (anonymous):

yes, but in this case i also i have write a algebraic expression and im not sure what i would write about this one.

jimthompson5910 (jim_thompson5910):

when we say "a number" we use a letter to represent an unknown number usually x is the letter we use x is just a placeholder. Think of it as a box where the number would go

jimthompson5910 (jim_thompson5910):

"a number times 2" means x times 2 = x*2 = 2*x = 2x

jimthompson5910 (jim_thompson5910):

with me so far?

OpenStudy (anonymous):

yea, i mean i know all that but i just dont know what order to put the phrase in, that;s what i need help with

OpenStudy (anonymous):

i mean in a expression

jimthompson5910 (jim_thompson5910):

what does "less" mean in this context?

OpenStudy (anonymous):

subtraction

jimthompson5910 (jim_thompson5910):

that is correct, so when we say "12 less" we mean "subtract off 12"

jimthompson5910 (jim_thompson5910):

so we have 2x and then we subtract off 12 to get 2x - 12

jimthompson5910 (jim_thompson5910):

"12 less than a number times 2" translates to \(\Large 2x - 12\) where x is the unknown number

OpenStudy (anonymous):

ty